Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
Allocca, Chiummo - 16:42 Tuesday 07 May 2019 (45832) Print this report
Considerations about the range evolution

Some considerations about the range evolution of the last weeks.

In figure 1, the time series of many useful signals is shown. This long period can be split in two parts: 

- range drop starting from April 22nd, lasting up until April 29th, when the TCS action was performed (figure 2)

During this period the sidebands power in the recycling cavity has an opposite trend with respect to the range, while the carrier recycling gain stays the same. It could be that this sideband power variation was triggered by the misfunctioning of the WI CO2 laser, although it's not expected that it should increase rather than decrease.

The carrier power increases on B1p: although this signal gets a significant contribution from the sidebands power, an increase of the carrier power is visible also on the B1p phase camera, which is only sensitive to the carrier contribution. However, the power on B2 doesn't vary accordingly, but seems to be rather constant. The hypothesis could be that the increase of carrier losses as seen by B1p (which would bring the B2 power to decrease) is compensated by a mismatch on one of the two cavities (which instead would bring to a B2 power increase). Moreover, the power on B8 decreases, pointing to a mismatch of the West arm.

The drop of the optical gain is probably related to the power decrease in the West arm: indeed, since DARM is locked, the DARM offset authomatically adjusts in order to keep the B1 power constant, and this could be the cause of the OG variation.

In order to check if the arms power variation could be triggered by the etalon, the DRUM 2 frequency has been monitored (WI drum mode). The 0.07 Hz variation (which corresponds to a temperature variation of about 0.08K), which doesn't explain the 2% of of West arm power decrease. 

- range increase after the TCS action, which decreases again after touching a peak on April 30th (figure 3) 

The tuning of the TCS has started a trend reversal of the range. The peak of sensitivity happens right after the maintenance which lasted about 8 hours. At this moment, the power on B1p is lower, sidebands are low, B2 is low (which would hint to reduced mismatch), while the power coupled to the West arm is higher, and the same for the optical gain. However, after the peak sensitivity period lasting about 6 hours, the worsening starts again. 

----------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

No conclusions as yet. There are hints suggesting that the matching towards the West arm is worsened, reducing the power coupled to the cavity - even though the cause of this behavior is not understood. This can trigger a variation of the optical gain, since the potential mismatch between the two reflected beams could induce a worse coupling to the OMCs and therefore less power on B1. 

 

Images attached to this report
Search
×

Warning

×